OpenStudy (anonymous):

What step am i missing? I don't understand how I can get to 1/5 * 2/3 (x+1)^(3/2) integrated from 3 to 8 from 1/5 * integral from 3 to 8 of (x+1)^(1/2) dx...

OpenStudy (anonymous):

\[1/5 \int\limits_{3}^{8}(x+1)^(1/2) \]

OpenStudy (anonymous):

to

OpenStudy (anonymous):

\[1/5 * 2/3 (x+1) ^ (3/2) integrated from 3 \to 8\]

OpenStudy (anonymous):

What's happening here is an application of the rule for integration that deals with functions raised to powers. The general version is: \[\int\limits_{a}^{b}x^n = \frac{ 1 }{ n+1 }x^{n+1}\] from a to b. In this case, n is 1/2, so plugging that value into the general form yields: \[(1/5)\int\limits_{3}^{8}x^{1/2}=(1/5)\frac{ 1 }{ 1/2 + 1 }x^{1/2 +1}=(1/5)\frac{ 1 }{ 3/2 }x^{3/2}=(1/5)\frac{ 2 }{ 3 }x^{3/2}\] from 3 to 8.

OpenStudy (anonymous):

but this is (x+1) to the 1/2 power. doesn't that change it?

OpenStudy (anonymous):

Ah, sorry I missed that. It doesn't change much about the answer, just replace x with (x + 1). There are cases where the difference would change the answer more dramatically, but since x and x+1 have the same derivative, there's no difference here.

OpenStudy (anonymous):

ok. thank you. just wondering, could that problem with different derivatives be solved by using the substitution method?

OpenStudy (anonymous):

Yeah, that would address that issue perfectly.

OpenStudy (anonymous):

thank you again for your help.